# Curiosity-Driven Development of Action and Language in Robots Through Self-Exploration

### v4

@@ Discussion
This study investigated how robots can develop action and language through self-exploration by integrating active inference with reinforcement learning. The experiments were designed to test four specific hypotheses, and the results provide clear support for each.
H1: Generalization is enhanced by compositional scale. The experiments confirmed that larger vocabularies of verbs, adjectives, and nouns led to greater generalization. Robots trained with richer compositional repertoires achieved higher success rates on unlearned actions, whereas smaller vocabularies constrained generalization severely. Our previous study (?) on supervised training of language and action for an arm robot also showed that compositional generalization improved as the size of verb–noun combinations in training increased. However, that work was limited in scale, examining only cases from $3\times 3$ to $5\times 8$ verb–noun combinations, where success rates for unlearned goals improved only from 57% to 71% under the condition of 80% training. By contrast, the present study examined much broader scaling, ranging from 48 to 180 possible compositions, while using only 33% of them for training. Under these conditions, generalization performance improved dramatically from near 25% to 85%. This contrast highlights that scale plays a critical role in enhancing generalization, and that curiosity-driven developmental learning provides a more powerful mechanism than supervised schemes under conditions of limited input. The finding also connects to the classical “poverty of the stimulus” problem raised by Chomsky (?), as it shows how compositionality enables powerful generalization from sparse training data. Once, we hypothesize that necessary training size could be proportional to summation of number of words appeared for each dimension instead of multiplication of it for all dimensions if compositionality size increases (?). This hypothesis becomes more plausible by the results in the current study which, however, should be confirmed in much more scaled experiments in the future.
@@
H5: Exception-handling rules exhibit U-shaped development. The results from Experiment 3 provide strong support for this hypothesis. When trained with two swapped command–action mappings, most robots initially produced the correct exceptional actions, subsequently regressed due to overgeneralization, and finally recovered the appropriate exception-specific mappings. This non-monotonic, U-shaped performance trajectory mirrors a well-established phenomenon in developmental psychology, in which children first succeed on irregular forms, later overgeneralize newly learned rules (e.g., producing “goed” for “went”), and ultimately reorganize their internal representations to master both rules and exceptions. Classic accounts interpret these dynamics as evidence for representational redescription—a restructuring of internal knowledge that enables more abstract, generative representations (?).
@@
Taken together, these findings demonstrate that curiosity-driven exploration, motor entropy, hierarchical acquisition of actions, and scalable compositional exposure jointly support efficient developmental learning of language and action. The parallels with infant development—rote-to-generalization progression, prerequisite learning, the role of vocabulary scale, and representation redescription—suggest that the mechanisms implemented here capture essential aspects of developmental psychology. More broadly, these results strengthen the view that reconstructing developmental processes in robots can offer insights into the “poverty of the stimulus” problem, showing how powerful generalization can arise from limited input when guided by intrinsic motivation, structured experience, and the principles of predictive coding and active inference (?, ?, ?, ?).
The current study is still limited in many aspects, and several possible extensions can be envisioned. One crucial limitation is that our experiments examined only a one-directional communication pathway from tutors to robots, relying on the command and feedback voices to guide the development of language–action mappings. In contrast, natural human development is characterized by interactive and bidirectional communication, where infants not only receive instructions but also actively solicit guidance, clarification, and scaffolding from caregivers.
@@
Another promising future direction concerns the development of robot–robot communication through the evolution of language. Previous research has explored this possibility from different perspectives: Steels introduced the framework of “language games” to study the emergence of shared vocabularies among agents (?), Miikkulainen and colleagues investigated the evolution of artificial language through evolutionary reinforcement learning (?), and Taniguchi proposed the emergence of symbols using a collective predictive coding approach (?). While these studies have demonstrated the possibility of emergent communication in an impressive manner, they still remain limited in that they mainly achieved the emergence of object labeling or naming, whereas the evolution of action-related language, such as verbs, has been much less explored. In this context, the current study based on active inference could be extended to address the evolution of dynamic linguistic structures, including verbs. Since our model implements active inference within a variational recurrent neural network, it is naturally suited for capturing temporal and dynamic aspects of action and language. A future extension of this work toward multi-robot interaction under the framework of “collective active inference” may thus provide novel insights into the evolution of embodied language, moving beyond static object labeling toward dynamic and action-oriented communication.
@@
$G_{t}$ $=-\underbrace{\eta D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{Curiosity/complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\alpha\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (1)
@@
The robot and the objects were simulated in PyBullet, the python physics simulator. Each wheel’s velocity was bounded within the range of [10,10][-10,10] meters per second. For scale, the robot’s body is a cube measuring 2 meters along each dimension (length, width, and height). The robot’s arm features two joints: yaw, which rotates left and right within a range of $30^{\degree},30^{\degree}]$, and pitch, which rotates forward and upward within a range of $0^{\degree},90^{\degree}].$ For smooth movement, the robot’s wheel and arm velocities were implemented with linear interpolation from the current to the target velocities.
We defined success criteria for each action category, which determined whether or not the robot earned an extrinsic reward. The distance between the robot and an object was measured from the object’s center to the center of the robot’s body. The robot was considered to be “facing the object” when the angular deviation between the robot’s forward direction the line connection it to the object was less than 15 degrees.
@@
Push Forward: The robot pushes the object farther than .1 meters, with respect to the robot’s facing direction. This must be maintained for 3 steps in a row.
Push Left: The robot pushes the object to the robot’s left farther than .2 meters, while the robot’s wheels have velocities below 5 meters per second (requiring use of the arm). This must be maintained for 3 steps in a row.
@@
There are some constraints in rewarding for actions which are described in the “Constraints in Performing Actions” subsection in the Supplementary Materials.
Developmental learning with Exception handling rules
To examine characteristics in developmental learning with exception handling rules, we tested an alternative training setup with a pair of swapped goal-action mappings. Robots were trained such that the command “watch magenta pillar” was only rewarded if the robot performed “be near green pole” instead, and vice versa. These exceptions conflicted with the regular compositional mappings. A matched control group of identical robots was trained without such exceptions. The success-rates of these robots is depicted in Fig. 7.
@@ Acknowledgments
Thank you to the Okinawa Institute of Science and Technology (OIST) and colleagues in OIST’s Cognitive Neurorobotics Unit and Neural Computation Unit for supporting this work. We also thank Taro Toyoizumi for valuable discussions on acquiring exception handling rules.
@@
For future reference, table S1 includes definitions for all relevant variables.
@@
In summary, the forward model is trained to minimize the evidence free energy $F$ (Eq. S3) by accurately reconstructing sensory observations and minimizing posterior complexity based on past experiences. Meanwhile, the actor-critic pair is trained to minimize expected free energy $G$, which includes an inverted complexity term (i.e., curiosity) and motor entropy to encourage exploration. This leads to emergent tension in an adversarial relationship: the actor is encouraged to maximize information gain by increasing the KL divergence between prior and posterior, which the forward model trains to minimize that same term. This establishes a dynamic push-pull effect, driving self-organized exploration. Please note that the imitation term in Eq. Free energy principle, Active Inference, and Kawahara Model depends on external demonstrations or expert policies; this term is ignored in our study, which focuses on self-exploration.
@@
    If the episode terminates at this step, the episode’s data is saved in a recurrent replay buffer. A batch of information is sampled from the buffer to train the forward model, actor, and critic. See the Supplementary Materials for more details.
@@
The second part of the sensory observation is the tactile sensation of touch. This is represented by a one value between 0 and 1 for each of the 16 sensors. Each value is equal to the fraction of time in the previous step during which the respective sensor was in contact with an object. See Fig. S3.
@@
The fourth and fifth parts of the sensation are the command voice and the tutor-feedback voices, which were described briefly in the Results section. Both voices are sequences of one-hot vectors. Table S5 displays the 18 terms (including silence) and their indexes in the one-hot vectors. For example, the command “Watch the Red Pillar” is represented by
@@ 17 Hourglass
In each step, the robot can only perform one of the six actions. This is implemented using definitions of actions and action prioritization. The actions Watch, Be Near, and Touch the Top cannot be performed simultaneously because of requirements regarding distance from the object and touching the object. The actions Push Left and Push Right cannot be performed simultaneously because of the directions of movements. If the robot satisfies the requirements for Touch the Top, we reject the actions Push Forward, Push Left, or Push Right. If the robot is performing Push Forward and Push Left or Push Right, we allow only the actions with the greatest distance pushed.
@@
10 robots trained in each way described in the Results section: no curiosity, sensory-motor curiosity, and all curiosity. The robots trained for 60000 epochs. In each epoch, the robot performed one episode which was saved in its recurrent replay buffer. Then the robot trained with a batch of 32 of its saved episodes.
@@
We measured the success-rates of these three types of robots in the six types of actions. The plots in Fig. 2 show the rolling average of success-rates of the three types robots from the beginning of training to the end of training after 60000 epochs, with 99% confidence intervals. Specifically, the plots show results of the robots regarding the goals with combinations of verb, adjective, and noun which the robots were not shown in training, testing for the ability to generalize vocabulary and syntax to unlearned combinations.
As we predicted in hypothesis $i$, robots with no curiosity performed the worst, with approximately 25% success-rate; robots with curiosity for sensory-motor observations performed better, with approximately 75% success-rate; and robots with curiosity for sensory-motor observations and the feedback voice are the best, with approximately 90% success-rate. As we predicted in hypothesis $ii$, the robot’s ability to perform simpler actions develop earliest, and the robot’s ability to perform more complex actions develop later, having required the simpler actions as prerequisites. Merely watching the object appears to be the simplest, developing earliest, while pushing object the object to the left or right appear to be the most complex, developing later.
@@
Experiment 2 tests the relationship between success-rates with learned goals and unlearned goals, specifically by robots using all curiosity. See Fig. 6. The left column shows success-rate plots of robots with learned actions, while the right column shows success-rate plots of robots with unlearned actions. The first row shows results for robots using the complete vocabulary: 6 verbs, 6 adjectives, and 5 nouns. The second and third row show results for robots trained with smaller vocabularies. In each of the three situations, the robots are trained with one third of the possible goals, and tested with the other two thirds.
@@
Movie S1. Example of training. Compares robot with curiosity mid-training and after training.
@@
To quantify U-shaped learning in exception handling, we scored the U-shaped structure of success-rate’s trajectories identifying non-monotonic developmental patterns consistent with representational redescription (?). The method combines robust smoothing, normalized scaling, and piecewise isotonic regression to fit a two-phase model with a central valley.

### v5

@@ Discussion
This study investigated how robots can develop action and language through self-exploration by integrating active inference with reinforcement learning. The experiments were designed to test five specific hypotheses, and the results provide clear support for each.
H1: Generalization is enhanced by compositional scale. The experiments confirmed that larger vocabularies of verbs, adjectives, and nouns led to greater generalization. Robots trained with richer compositional repertoires achieved higher success rates on unlearned actions, whereas smaller vocabularies constrained generalization severely. Our previous study (?) on supervised training of language and action for an arm robot also showed that compositional generalization improved as the size of verb-noun combinations in training increased. However, that work was limited in scale, examining only cases from $3\times 3$ to $5\times 8$ verb-noun combinations, where success rates for unlearned goals improved only from 57% to 71% under the condition of 80% training. By contrast, the present study examined much broader scaling, ranging from 48 to 180 possible compositions, while using only 33% of them for training. Under these conditions, generalization performance improved dramatically from near 25% to 85%. This contrast highlights that scale plays a critical role in enhancing generalization, and that curiosity-driven developmental learning provides a more powerful mechanism than supervised schemes under conditions of limited input. The finding also connects to the classical “poverty of the stimulus” problem raised by Chomsky (?), as it shows how compositionality enables powerful generalization from sparse training data. Once, we hypothesized that necessary training size could be proportional to summation of number of words appeared for each dimension instead of multiplication of it for all dimensions if compositionality size increases (?). This hypothesis becomes more plausible given the results in this current study, which should be confirmed in much more scaled experiments in the future.
@@
H5: Exception-handling rules exhibit U-shaped development. The results from Experiment 3 provide strong support for this hypothesis. When robots were trained with two swapped command–action mappings, U-shaped performance trajectories were observed more frequently than robots trained without these exceptions. Our analysis of the latent variables in the model network showed that overgeneralization takes place in the middle of development and such internal representation is redescribed to accommodate with exception-handling rules later. This non-monotonic, U-shaped performance trajectory mirrors a well-established phenomenon in developmental psychology, in which children first succeed on irregular forms, later overgeneralize newly learned rules (e.g., producing “goed” for “went”), and ultimately reorganize their internal representations to master both rules and exceptions. Classic accounts interpret these dynamics as evidence for representational redescription, a restructuring of internal knowledge that enables more abstract, generative representations (?).
@@
Taken together, these findings demonstrate that curiosity-driven exploration, motor entropy, hierarchical acquisition of actions, and scalable compositional exposure jointly support efficient developmental learning of language and action. The parallels with infant development, including rote-to-generalization progression, prerequisite learning, the role of vocabulary scale, and representation redescription, suggest that the mechanisms implemented here capture essential aspects of developmental psychology. More broadly, these results strengthen the view that reconstructing developmental processes in robots can offer insights into the “poverty of the stimulus” problem, showing how powerful generalization can arise from limited input when guided by intrinsic motivation, structured experience, and the principles of predictive coding and active inference (?, ?, ?, ?).
The current study is still limited in many aspects, and several possible extensions can be envisioned. One crucial limitation is that our experiments examined only a one-directional communication pathway from tutors to robots, relying on the command and feedback voices to guide the development of language-action mappings. In contrast, natural human development is characterized by interactive and bidirectional communication, where infants not only receive instructions but also actively solicit guidance, clarification, and scaffolding from caregivers.
@@
Another promising future direction concerns the development of robot-robot communication through the evolution of language. Previous research has explored this possibility from different perspectives: Steels introduced the framework of “language games” to study the emergence of shared vocabularies among agents (?), Miikkulainen and colleagues investigated the evolution of artificial language through evolutionary reinforcement learning (?), and Taniguchi proposed the emergence of symbols using a collective predictive coding approach (?). While these studies have demonstrated the possibility of emergent communication in an impressive manner, they still remain limited in that they mainly achieved the emergence of object labeling or naming, whereas the evolution of action-related language, such as verbs, has been much less explored. In this context, the current study based on active inference could be extended to address the evolution of dynamic linguistic structures, including verbs. Since our model implements active inference within a variational recurrent neural network, it is naturally suited for capturing temporal and dynamic aspects of action and language. A future extension of this work toward multi-robot interaction under the framework of “collective active inference” may thus provide novel insights into the evolution of embodied language, moving beyond static object labeling toward dynamic and action-oriented communication.
@@
$G_{t}$ $=-\underbrace{\eta D_{KL}[q(z_{t}|o_{t},a_{t-1},h_{t-1})||p(z_{t}|a_{t-1},h_{t-1})]}_{\text{Curiosity/complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\alpha\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (3)
@@
The robot and the objects were simulated in PyBullet, the python physics simulator. Each wheel’s velocity was bounded within the range of [10,10][-10,10] meters per second. For scale, the robot’s body is a cube measuring 2 meters along each dimension (length, width, and height). The robot’s arm features two joints: yaw, which rotates left or right within a range of $30^{\degree},30^{\degree}]$, and pitch, which rotates forward or upward within a range of $0^{\degree},90^{\degree}].$ For smooth movement, the robot’s wheel and arm velocities were implemented with linear interpolation from the current to the target velocities.
We defined success criteria for each action category, which determined whether or not the robot earned an extrinsic reward by completing a goal. The distance between the robot and an object was measured from the object’s center to the center of the robot’s body. The robot was considered to be “facing the object” when the angular deviation between the robot’s forward direction and the line connection it to the object was less than 15 degrees.
@@
Push Forward: The robot pushes the object farther than .1 meters with respect to the robot’s facing direction. This must be maintained for 3 steps in a row.
Push Left: The robot pushes the object to the robot’s left farther than .2 meters while the robot’s wheels have velocities below 5 meters per second (requiring use of the arm). This must be maintained for 3 steps in a row.
@@
There are constraints in rewarding for actions which are described in the “Constraints in Performing Actions” subsection in the Supplementary Materials.
@@ Acknowledgments
Thank you to the Okinawa Institute of Science and Technology (OIST) and colleagues in OIST’s Cognitive Neurorobotics Unit and Neural Computation Unit for supporting this work. We also thank Taro Toyoizumi for valuable discussions on acquiring exception-handling rules.
@@
For future reference, table S1 includes definitions for relevant variables.
@@
In summary, the forward model is trained to minimize the evidence free energy $F$ (Eq. S3) by accurately reconstructing sensory observations and minimizing posterior complexity based on past experiences. Meanwhile, the actor-critic pair is trained to minimize expected free energy $G$, which includes an inverted complexity term (i.e., curiosity) and motor entropy to encourage exploration. This leads to emergent tension in an adversarial relationship: the actor is encouraged to maximize information gain by increasing the KL divergence between prior and posterior, while the forward model trains to minimize that same term. This establishes a dynamic push-pull effect, driving self-organized exploration. Please note that the imitation term in Eq. Free energy principle, Active Inference, and Kawahara Model depends on external demonstrations or expert policies; this term is ignored in our study, which focuses on self-exploration.
@@
    If the episode terminates at this step, the episode’s data is saved in a recurrent replay buffer. A batch of information is sampled from the buffer to train the forward model, actor, and critic.
@@
The second part of the sensory observation is the tactile sensation of touch. This is represented by one value between 0 and 1 for each of the robot’s 16 sensors. Each value is equal to the fraction of time in the previous step during which the respective sensor was in contact with an object. See Fig. S3.
@@
The fourth and fifth parts of the sensation are the command voice and the tutor-feedback voice, which were described briefly in the Results section. Both voices are sequences of one-hot vectors. Table S5 displays the 18 words (including silence) and their indexes in the one-hot vectors. For example, the command “Watch the Red Pillar” is represented by
@@ 17 Hourglass
In each step, the robot can only perform one of the six actions. This is implemented using definitions of actions and action prioritization. The actions Watch, Be Near, and Touch the Top cannot be performed simultaneously because of requirements regarding distance from the object and touching the object. The actions Push Left and Push Right cannot be performed simultaneously because of the directions of movements. If the robot satisfies the requirements for Touch the Top, we reject the actions Push Forward, Push Left, or Push Right. If the robot is performing Push Forward and Push Left or Push Right, we accept only the action with the greatest distance pushed.
@@
10 robots are trained in each way described in the Results section: no curiosity, sensory-motor curiosity, and all curiosity. The robots trained for 60000 epochs. In each epoch, the robot performed one episode which was saved in its recurrent replay buffer. Then the robot trained with a batch of 32 of its saved episodes.
@@
We measured the success-rates of these three types of robots in the six types of actions. The plots in Fig. 3 show the rolling average of success-rates of the three types of robots from the beginning of training to the end of training after 60000 epochs, with 99% confidence intervals. Specifically, the plots show results of the robots regarding the goals with combinations of verb, adjective, and noun which the robots were not shown in training, testing for the ability to generalize vocabulary and syntax to unlearned combinations.
As we predicted in hypothesis $i$, robots with no curiosity performed the worst, with approximately 25% success-rate; robots with sensory-motor curiosity performed better, with approximately 75% success-rate; and robots with all curiosity are the best, with approximately 90% success-rate. As we predicted in hypothesis $ii$, the robot’s ability to perform simpler actions develop earliest, and the robot’s ability to perform more complex actions develop later, having required the simpler actions as prerequisites. Merely watching the object appears to be the simplest, developing earliest, while pushing object the object to the left or right appear to be the most complex, developing later.
Fig. S5 shows PCA results for the estimated posterior latent states of one robot which was not supported with curiosity. In contrast to Fig. 4, this robot has heavily entangled understandings of some verbs.
Fig. S5: PCA for language latent variables in the case of no curiosity. PCA applied to latent representations of command voice inputs after complete development.
@@
Experiment 2 tests the relationship between success-rates with learned goals and unlearned goals, specifically by robots using all curiosity. See Fig. 6. The left column shows robot’s success-rates with learned actions, while the right column shows robot’s success-rates with unlearned actions. The first row shows results for robots using the complete vocabulary: 6 verbs, 6 adjectives, and 5 nouns. The second and third row show results for robots trained with smaller vocabularies. In each of the three situations, the robots are trained with one third of the possible goals, and tested with the other two thirds.
@@
Movie S1. Example of training. Compares a robot with all curiosity mid-training and after training.
@@
To quantify U-shaped learning in exception-handling, we scored the U-shaped structure of success-rate trajectories identifying non-monotonic developmental patterns consistent with representational redescription (?). The method combines robust smoothing, normalized scaling, and piecewise isotonic regression to fit a two-phase model with a central valley.